(6 points) Use the Laplace transform to solve the following initial value problem: y + 3y = 0 y(0) = -3, y(0) = 6 First, u
Answer #1

y + 3y = 0. taking Laplace drains finalin s yrs) - sy 109 - yro2+ 384()– 3410930. (s?+38)y +35–6+9=0. (s+ +38) Y + 35+3 >